# Vision-Based Mistake Analysis in Procedural Activities: A Review of Advances and Challenges

### v1

@@ VII-B1 Ordinary Classification (OC)
The methods [19], [21], [13] following this learning strategy treat both action and mistake recognition (and detection) as standard supervised classification tasks. A backbone network, such as a 3D convolutional model (e.g., I3D [32]) or a transformer-based architecture (e.g., TimeSformer [48]), is trained end-to-end using a multi-class cross-entropy loss for classifying action steps and a binary cross-entropy loss for labeling each step as correct or erroneous. Crucially, this approach requires dense, segment-level annotations, including both the ground-truth action class and an associated correctness label for each step. As such, it necessitates the availability of annotated sequences containing both correct and erroneous executions during training, which can be costly and time-intensive to curate, particularly in domains with complex or long-horizon tasks.
@@ VII-B2 One-Class Classification (OCC)
In this context, models for mistake recognition or detection are trained solely on samples of correct behavior, aiming to identify deviations from the learned norm during inference. This approach suits mistake and broader anomaly detection tasks, where anomalous events, such as procedural mistakes, are rare and lack explicit error examples for training. Existing works follow different strategies within this framework. Some adopt prototype learning, clustering normal executions as reference points at test time. For instance, Lee et al. [15] learn multiple contrastive prototypes per procedural step and score incoming frames by their distance to the nearest prototype, while Seminara et al. [45] extend this idea using a differentiable task graph that models step transitions and flags off-graph deviations as errors.
Other approaches explore OCC with predictive and step-aware reasoning strategies, in which the model anticipates future procedural steps and evaluates the consistency of observed behavior against these predictions. Methods such as PREGO (in Flaborea et al. [24]) and the extension TI-PREGO (in Plini et al. [37]) utilize an OCC framework that anticipates the next step embedding based on symbolic task progression predicted by an LLM, while grounding the actual observation through a visual encoder. Discrepancies between predicted and actual embeddings are used to detect errors as they occur in real time, without ever observing mistakes during training.
@@ VII-B3 Task Completion (TC)
Training objectives in this category often include masked step prediction, sequence reconstruction, or permutation-based reasoning, encouraging the model to infer and enforce task-level consistency. Unlike conventional classification frameworks, TC approaches do not require explicit error labels; correctness is inferred from whether the predicted trajectory conforms to a learned procedural schema. For instance, Narasimhan et al. [51] learn structured task representations by predicting randomly masked steps in instructional videos, allowing the model to identify steps that break the expected temporal or logical flow. In a complementary formulation, a subset of methods approach this objective via anticipation-based reasoning, where correctness is evaluated by comparing the model’s recognized step with the expected next step, as inferred from a structured procedure. PREGO [24] and TI-PREGO [37] formulate mistake detection as a joint problem of step recognition and step anticipation, where a mismatch between anticipated and recognized steps signals a deviation from the canonical trajectory of the action-steps of the activity.
Method Task Error Error Error Dataset View Learning Modality Superv. Procedural Online Video Action & Activity
@@
Moriwaki et al. [19] MD PEs, Coarse TALM, ASA BRIO-TA [19] Ego OC RGB Fully STM I3D [32] + MS-TCN [35]
ICIP’22 $E$s
Ding et al. [16] MD PEs Coarse, CPM Assembly101 [14] NV TC Text (AL) Full STM, SKG TempAgg [54]
@@
Narasimhan et al. [51] MD PEs Coarse CPM COIN [55] Ego/ TC RGB Weak TDM VideoTaskformer [51]
@@
Wang et al. [21] MD PEs, Coarse CPM HoloAssist [21] Ego OC RGB, Full STM TimeSformer [48]
@@
Ghoddoosian et al. [20] MD PEs Coarse CPM, TALM ATA [20], CSV [18] Exo OC, OCC, RGB Weak TDM I3D [32] + OadTR [56]
ICCV’23 VTA + Viterbi variant [20]
Schoonbeek et al. [23] MR PEs Coarse POS, F1 IndustReal [23] Ego TC, OCC RGB, VL, Weak TDM Both MViTv2 [57]
WACV’24 Stereo
Peddi et al. [13] MTR, PEs, Coarse, CPM CaptainCook4D [13] Ego $C$ RGB Full SFM (MTR, Omnivore [40]
NeurIPS’24 EMP, $E$s Fine EMP), STM (D) + ActionFormer [34]
@@
Lee et al. [15] MD PEs Coarse EDA, AUC, EgoPER [15], ATA [20] Ego OCC RGB Full SFM (I3D[32]+FasterRCNN [58]
CVPR’24 O-IoU, TALM HoloAssist [21] +GCN) + ActionFormer [34]
Flaborea et al. [24] MD PEs Coarse CPM Assembly101-O [24], Ego TC, OCC RGB Weak TDM Both OadTR [56] + Llama 2.0
CVPR’24 Epic-Tent-O [24]
Seminara et al. [45] MD PEs Coarse CPM Assembly101-O [24], Ego TC, OCC RGB Weak PTG Both MinoRoad [36]
NeurIPS’24 Epic-Tent-O [24] + Task Graph Transformer [45]
Haneji et al. [12] MD PEs Coarse TALM EgoOops [12] Ego VTA, TC RGB Fully TDM, STM StepFormer [59]
ArXiv’24 + Drop-DTW [60]
Plini et al. [37] MD PEs Coarse CPM Assembly101-O [24], Ego TC, OCC RGB Weak TDM Both MiniRoad [36]
ArXiv’24 Epic-Tent-O [24] + Llama 3.1 (CoT)
Mazzamuto et al. [46] MD $E$s Coarse OCC HoloAssist [21], Ego SFM RGB, EG Both Gaze-centered Transformer
CVPR’25 Epic-Tent [17] Autoenc. + MoCoDAD [61]
Kung et al. [26] MD PEs, Coarse EDA, AUC EgoPER [15] Ego VTA, TC RGB, Weak PTG Llama 3 + ASFormer [62]
ArXiv’25 $E$s Text + HierVL [63]
TABLE II: Method taxonomy across key design factors. L.: learning. Tasks: MR–mistake recognition, MTR–mistake type recognition, EMP–early mistake recognition, MD–mistake detection. NV: non-vision method. Modalities: AL–action labels, HP–hand pose, EG–eye gaze, VL–visible light. Procedural structure: STM–stepwise model, TDM–template-driven model, SKG–semantic knowledge graph, PTG–procedural task graph, SF–structure-free. Error metrics: CPM–classification metrics (Acc, Prec, Rec, F1, AUC), TALM–temporal localization metrics (Edit Distance, IoU, F1@0.5@0.5), EDA–error detection accuracy, POS–procedure order similarity, ASA–anomaly section accuracy, OIoU–omission IoU. Misc.: GCN–graph convolutional networks, ACoT–Automatic Chain of Thought [64]. For works introducing novel datasets, only the top-performing baseline is listed in the Method Overview column.
@@ VII-B4 Video-Text Alignment as a Supervisory Signal (VTA)
VTA-based approaches utilize procedural text (e.g., step descriptions from manuals, task guides, or LLM-generated scripts) as a supervisory signal to guide action detection, anticipation, and activity understanding [65, 66, 67]. These methods typically segment untrimmed videos into candidate action clips by aligning them to textual templates using techniques such as dynamic time warping, cross-modal attention, or contrastive learning between video and language embeddings, then pass the aligned segments to classification or matching modules to predict the activity or next action.
In mistake analysis, VTA enables assessing step correctness by evaluating alignment strength or semantic consistency between predicted visual content and the textual description of correct behavior. Its adoption remains limited due to scarce datasets like EgoOops [12] and CaptainCook4D [13], which provide paired instructional text and annotated mistakes. In EgoOops, Haneji et al. [12] align egocentric video segments with procedural transcripts. StepFormer++ predicts the current step and localizes mistakes by detecting misalignments between visual input and expected procedural steps, using cross-attention to learn temporal and semantic alignment between video tokens and step embeddings. To address data limitations, Kung et al. [26] propose a counterfactual VTA framework for mistake analysis using Ego4D [22]. While Ego4D lacks error annotations, it provides rich activity- and action-level captions. LLaMA 3.2 8B is used to generate counterfactual descriptions at action and activity levels, describing missing, misordered, or incorrectly executed steps. These synthetic error-augmented captions are paired with correct video clips for contrastive training of HierVL [63], a hierarchical video-language model.
@@ VII-C1 Fully Supervised Approaches
Fully supervised approaches rely on datasets with explicit error annotations at a fine-grained level. In this setting, mistakes are annotated with precise spatial and temporal localization, typically at the frame or segment level, enabling models to learn the error type and also its exact temporal occurrence in the video. Most methods following this paradigm operate in an offline setting [13], [14], [21], where detailed annotations are used to supervise both step segmentation and mistake classification. Ding et al. [16] leverage fine-grained labels to construct spatiotemporal knowledge graphs guiding mistake reasoning. Similarly, Mirowaki et al. [19], in the BRIO-TA dataset, provide segment-level annotations indicating correct or erroneous execution, allowing models to be trained using standard supervised classification objectives at the action segment level.
@@ VII-C2 Weakly-Supervised Methods
In mistake analysis, the first perspective concerns the absence of explicit mistake annotations, compensated by exploiting the procedural structure (actions and their transitions) of the activity. This is exemplified by methods leveraging global step order to detect execution inconsistencies [25, 20, 23, 51]. Although these methods do not require annotated mistake labels for training, they still rely on step-level annotations, i.e. knowledge of actions and their order. The second perspective relates to annotation granularity, where weak supervision labels mistakes only at the video level (e.g., indicating a mistake’s presence) without specifying when it occurs. In such cases, models must implicitly localize deviations based on high-level error cues (e.g., “action skipped”).
@@
Grauman et al. in Ego-Exo4D [25] explore a weak mistake detection strategy with two variants depending on supervision granularity: (1) instance-level supervision, where both video segments and key-step labels are provided during training and inference, analogous to action recognition, and (2) procedure-level supervision, where training and inference use unlabeled segments along with a taxonomy of procedure-specific key-step names. In the second approach, key-step assignments are generated via a pre-trained video-language model, providing pseudo-labels and exemplifying self-supervised weak supervision. Similarly, Ghoddoosian et al. [20] propose a weakly-supervised method for instructional mistake detection using only activity transcripts without frame-level labels or error examples. Supervision is limited to expected action sequences, with no temporal error information. Their method infers frame-wise action boundaries by aligning predictions to the transcript using a loss that enforces consistency with valid sequences while maximizing separation from invalid ones. Narasimhan et al. [51], in VideoTaskformer, operate under weak supervision differently: instead of aligning to a transcript, the model learns procedural knowledge by predicting masked steps within instructional sequences. Trained with step-level annotations (e.g., “crack egg”, “whisk”, “pour”), without step timing or correctness information, it infers missing steps based on surrounding context, acquiring a representation of procedural structure that identifies absent or out-of-order steps.
@@
Finally, a recent sub-line of weakly-supervised approaches [24, 37] departs from previous formulations and defines mistake detection as a combination of step recognition and step anticipation. These methods exploit the structured nature of task transcripts as supervision signals, without requiring any frame-level action or mistake annotations. Notably, they leverage pre-trained LLMs as reasoning agents to infer expected next steps based on procedural understanding.

### v2

@@ VII-B1 Ordinary Classification (OC)
The methods [BRIOTA], [HoloAssist], [CaptainCook4D], [patsch2025mistsense] following this learning strategy treat both action and mistake recognition (and detection) as standard supervised classification tasks. A backbone network, such as a 3D convolutional model (e.g., I3D [carreira2017quo]) or a transformer-based architecture (e.g., TimeSformer [bertasius2021space]), is trained end-to-end using a multi-class cross-entropy loss for classifying action steps and a binary cross-entropy loss for labeling each step as correct or erroneous. Crucially, this approach requires dense, segment-level annotations, including both ground-truth action class and associated correctness label for each step. As such, it necessitates the availability of annotated sequences containing both correct and erroneous executions during training, which can be costly and time-intensive to curate, particularly in domains with complex or long-horizon tasks.
@@ VII-B2 One-Class Classification (OCC)
In this context, models for mistake recognition or detection are trained solely on samples of correct behavior, aiming to identify deviations from the learned norm during inference. This approach suits mistake and broader anomaly detection tasks, where anomalous events, such as procedural mistakes, are rare and lack explicit error examples for training. Existing works follow different strategies within this framework. Some adopt prototype learning, clustering normal executions as reference points at test time. For instance, Lee et al. [EgoPED] learn multiple contrastive prototypes per procedural step and score incoming frames by their distance to the nearest prototype, while Seminara et al. [seminara2024differentiable] extend this idea using a differentiable task graph that models step transitions and flags off-graph deviations as errors. However, static representations struggle to capture the diversity of branching tasks. Addressing this, Huang et al. [huang2025modeling] introduce the AMNAR framework, which, similar to the graph-guided approach of Seminara et al. [seminara2024differentiable], leverages a task graph to model valid transitions. AMNAR uses the graph to dynamically reconstruct multiple valid action representations conditioned on the current context, ensuring that permissible variations in execution order are not misclassified as mistakes.
Other approaches explore OCC with predictive and step-aware reasoning strategies, in which the model anticipates future procedural steps and evaluates the consistency of observed behavior against these predictions. Methods such as PREGO (in Flaborea et al. [flaborea2024prego]) and the extension TI-PREGO (in Plini et al. [plini2024ti]) utilize an OCC framework that anticipates the next step embedding based on symbolic task progression predicted by an LLM, while grounding the actual observation through a visual encoder. Discrepancies between predicted and actual embeddings are used to detect errors as they occur in real time, without ever observing mistakes during training.
@@ VII-B3 Task Completion (TC)
Training objectives in this category often include masked step prediction, sequence reconstruction, or permutation-based reasoning, encouraging the model to infer and enforce task-level consistency. Unlike conventional classification frameworks, TC approaches do not require explicit error labels; correctness is inferred from whether the predicted trajectory conforms to a learned procedural schema. For instance, Narasimhan et al. [narasimhan2023learning] learn structured task representations by predicting randomly masked steps in instructional videos, allowing the model to identify steps that break the expected temporal or logical flow. In a complementary formulation, a subset of methods approach this objective via anticipation-based reasoning, where correctness is evaluated by comparing the model’s recognized step with the expected next step, as inferred from a structured procedure. PREGO [flaborea2024prego] and TI-PREGO [plini2024ti] formulate mistake detection as a joint problem of step recognition and step anticipation, where a mismatch between anticipated and recognized steps signals a deviation from the canonical trajectory of the action-steps of the activity.
Method Task Error Error Error Dataset View Learning Modality/Input Superv. Procedural Online Video Action & Activity
@@
Moriwaki et al. [BRIOTA] MD PEs, Coarse TALM, BRIO-TA [BRIOTA] Ego OC RGB Fully STM I3D [carreira2017quo] + MS-TCN [farha2019ms]
ICIP’22 $E$s ASA
Ding et al. [ding2023mistakecountsassembly] MD PEs Coarse, CPM Assembly101 [Assembly101] NV TC Text (AL) Full STM, SKG TempAgg [sener2020temporal]
@@
Narasimhan et al. [narasimhan2023learning] MD PEs Coarse CPM COIN [COIN] Ego/ TC RGB Weak TDM VideoTaskformer [narasimhan2023learning]
@@
Wang et al. [HoloAssist] MD PEs, Coarse CPM HoloAssist [HoloAssist] Ego OC RGB, Full STM TimeSformer [bertasius2021space]
@@
Ghoddoosian et al. [ATA] MD PEs Coarse CPM, ATA [ATA], CSV [CSV] Exo OC, OCC, RGB Weak TDM I3D [carreira2017quo] + OadTR [wang2021oadtr]
ICCV’23 TALM VTA + Viterbi variant [ATA]
Schoonbeek et al. [IndustReal] MR PEs Coarse POS IndustReal [IndustReal] Ego TC, OCC RGB, VL, Weak TDM Both MViTv2 [li2022mvitv2]
WACV’24 CPM (F1) Stereo
Peddi et al. [CaptainCook4D] MTR, PEs, Coarse, CPM CaptainCook4D [CaptainCook4D] Ego $C$ RGB Full SFM (MTR, Omnivore [girdhar2022omnivore]
NeurIPS’24 EMP, $E$s Fine EMP), STM (D) + ActionFormer [zhang2022actionformer]
@@
Lee et al. [EgoPED] MD PEs Coarse EDA, CPM (AUC), EgoPER [EgoPED], ATA [ATA] Ego OCC RGB Full SFM (I3D[carreira2017quo]+FasterRCNN [ren2016faster]
CVPR’24 O-IoU, TALM HoloAssist [HoloAssist] +GCN) + ActionFormer [zhang2022actionformer]
Flaborea et al. [flaborea2024prego] MD PEs Coarse CPM Assembly101-O [flaborea2024prego], Ego TC, OCC RGB Weak TDM Both OadTR [wang2021oadtr] + Llama 2.0
CVPR’24 Epic-Tent-O [flaborea2024prego]
Seminara et al. [seminara2024differentiable] MD PEs Coarse CPM Assembly101-O [flaborea2024prego], Ego TC, OCC RGB Weak PTG Both MinoRoad [an2023miniroad] +
NeurIPS’24 Epic-Tent-O [flaborea2024prego] + Task Graph Transformer [seminara2024differentiable]
Haneji et al. [EgoOops] MD PEs Coarse TALM EgoOops [EgoOops] Ego VTA, TC RGB Fully TDM, STM StepFormer [dvornik2023stepformer]
arXiv’24 + Drop-DTW [dvornik2021drop]
Plini et al. [plini2024ti] MD PEs Coarse CPM Assembly101-O [flaborea2024prego], Ego TC, OCC RGB Weak TDM Both MiniRoad [an2023miniroad]
arXiv’24 Epic-Tent-O [flaborea2024prego] + Llama 3.1 (CoT)
Mazzamuto et al. [mazzamuto2025gazing] MD $E$s Coarse CPM HoloAssist [HoloAssist], Ego OCC RGB, SFM Both Gaze-centered Transformer
CVPR’25 Epic-Tent [EPICTENT] EG Autoenc. + MoCoDAD [flaborea2023multimodal]
Huang et al. [huang2025modeling] MD PEs Coarse EDA, EgoPER [EgoPED], HoloAssist [HoloAssist], Ego OCC RGB, Weak PTG Both I3D [carreira2017quo] + ActionFormer [zhang2022actionformer]
CVPR’25 CPM (Prec, AUC) Captain-Cook4D [CaptainCook4D] TG +Graph Structure Operations
Hou et al. [ICME] MD $E$s Coarse EDA HoloAssist [HoloAssist], Ego OCC RGB Weak STM Both (ActionFormer [zhang2022actionformer]+ CDC)+
ICME’25 PEs** CPM (AUC) EgoPER [EgoPED] GMM+Gaussian Smoothing Module
Kung et al. [kung2025changed] MD PEs, Coarse EDA, EgoPER [EgoPED] Ego VTA, TC RGB, Weak PTG Llama 3 + ASFormer [yi2021asformer]
arXiv’25 $E$s CPM (AUC) Text + HierVL [ashutosh2023hiervl]
Lee and Elhamifar [lee2025error] MD PEs Coarse TALM, EgoPER [EgoPED], Ego TC RGB, Weak PTG Both DiffAct [liu2023diffusion] + G2Vid [dvornik2022flow]
ICCV’25 $E$s Fine CPM CaptainCook4D*[CaptainCook4D] TG + (VideoCLIP [Xu2021VideoCLIPCP] + GPT4o mini)
(ViT [dosovitskiy2020image] + Q-Former [li2023blip]) +
Patsch et al. [patsch2025mistsense] MD-e PEs, Coarse, CPM (F1) Epic-Tent-O [flaborea2024prego], Ego OC, RGB, Fully STM Both (HardFormer [shamil2024utility] + Q-Former [li2023blip])
ICCV’25 $E$s Fine NGM HoloAssist [HoloAssist] VTA HP + Llama 2.0
Li et al. [li2025mistake] (ST) MD $E$s Coarse CPM, TALM, EK-M [li2025mistake], Ego VTA RGB Fully STM SRL [gardner2018allennlp] + InternVideo2 [wang2024internvideo2]
arXiv’25 PEs** Fine OBJM Ego4D-M [li2025mistake] Text + Transformer Heads
TABLE II: Method taxonomy across key design factors. Tasks: {MR–mistake recognition, MTR–mistake type recognition, EMP–early mistake recognition, MD–mistake detection (temporal), (ST)MD - spatio-temporal MD, -e: supports explainability}. Modalities: {AL–action labels, HP–hand pose, EG–eye gaze, VL–visible light, TG-task graph}. Procedural structure: {STM–stepwise model, TDM–template-driven model, SKG–semantic knowledge graph, PTG–procedural task graph, SF–structure-free}. Error metrics: {CPM–classification metrics (Acc, Prec, Rec, F1, AUC), TALM–temporal localization metrics (Edit Distance, IoU, F1@0.5@0.5), EDA–error detection accuracy, POS–procedure order similarity, ASA–anomaly section accuracy, OIoU–omission IoU, NGM: n-gram metrics (BLEU, ROUGE-L, CIDEr).}. Misc.: {NV: non-vision method, OBJM-object detection metrics, GCN–graph convolutional networks, ACoT–Automatic Chain of Thought [zhang2022automatic], TAS-Temporal Action Segmentation, CDC - Causal Dilated Convolution, GMM- Gaussian Mixture Model}. Note that: For works introducing novel datasets, the top-performing baseline is listed in the last column. Overall, for every work we list top-performing model configuration. *: evaluated on a subset of the dataset.**:[ICME] does not explicitly model procedural structure; such errors are detected when manifested as deviations in the per-step visual distribution.
@@ VII-B4 Video-Text Alignment as a Supervisory Signal (VTA)
VTA-based approaches utilize procedural text (e.g., step descriptions from manuals, task guides, or LLM-generated scripts) as a supervisory signal to guide action detection, anticipation, and activity understanding [kahatapitiya2024victr, Chen_2024_CVPR, li2024topa]. These methods typically segment untrimmed videos into candidate action clips by aligning them to textual templates using techniques such as dynamic time warping, cross-modal attention, or contrastive learning between video and language embeddings, then pass the aligned segments to classification or matching modules to predict the activity or next action.
In mistake analysis, VTA enables assessing step correctness by evaluating alignment strength or semantic consistency between predicted visual content and the textual description of correct behavior. Its adoption remains limited due to scarce datasets like EgoOops [EgoOops] and CaptainCook4D [CaptainCook4D], which provide paired instructional text and annotated mistakes. In EgoOops, Haneji et al. [EgoOops] align egocentric video segments with procedural transcripts. StepFormer++ predicts the current step and localizes mistakes by detecting misalignments between visual input and expected procedural steps, using cross-attention to learn temporal and semantic alignment between video tokens and step embeddings. To address data limitations, Kung et al. [kung2025changed] propose a counterfactual VTA framework for mistake analysis using Ego4D [grauman2022ego4d]. While Ego4D lacks error annotations, it provides rich activity- and action-level captions. LLaMA 3.2 8B is used to generate counterfactual descriptions at action and activity levels, describing missing, misordered, or incorrectly executed steps. These synthetic error-augmented captions are paired with correct video clips for contrastive training of HierVL [ashutosh2023hiervl], a hierarchical video-language model.
Moving beyond sentence-level alignment, Li et al. [li2025mistake] in MisFormer, which utilizes Semantic Role Labeling (SRL) to decompose procedural instructions into fine-grained components (e.g., Predicate (verb), Object). Unlike previous VTA approaches that treat the text as a single embedding, MisFormer employs a dual-branch transformer that cross-attends these specific role tokens against video patches. This enables the model to determine specifically which part of the instruction was violated (e.g., correct action (verb) but wrong object). Complementing these attribution-focused methods, Patsch et al. [patsch2025mistsense] leverage VTA for post-hoc explanation rather than just detection. Their MistSense framework employs a gating mechanism that, upon detecting an error, projects video-based features into the embedding space of a Large Language Model (LLM). This allows the system to generate natural language descriptions of the error (e.g., “The person is mounting a screw in a wrong way”), bridging the gap between binary detection and interpretable user feedback.
@@ VII-C1 Fully Supervised Approaches
Fully supervised approaches rely on datasets with explicit error annotations at a fine-grained level. In this setting, mistakes are annotated with precise spatial and temporal localization, typically at the frame or segment level, enabling models to learn the error type and also its exact temporal occurrence in the video. Most methods following this paradigm operate in an offline setting [CaptainCook4D], [Assembly101], [HoloAssist], where detailed annotations are used to supervise both step segmentation and mistake classification. Ding et al. [ding2023mistakecountsassembly] leverage fine-grained labels to construct spatiotemporal knowledge graphs guiding mistake reasoning. Similarly, Mirowaki et al. [BRIOTA], in the BRIO-TA dataset, provide segment-level annotations indicating correct or erroneous execution, allowing models to be trained using standard supervised classification objectives at the action segment level.
@@ VII-C2 Weakly-Supervised Methods
In mistake analysis, the first perspective concerns the absence of explicit mistake annotations, compensated by exploiting the procedural structure (actions and their transitions) of the activity. This is exemplified by methods leveraging global step order to detect execution inconsistencies [grauman2024ego, ATA, IndustReal, narasimhan2023learning, lee2025error]. Although these methods do not require annotated mistake labels for training, they still rely on step-level annotations, i.e. knowledge of actions and their order. The second perspective relates to annotation granularity, where weak supervision labels mistakes only at the video level (e.g., indicating a mistake’s presence) without specifying when it occurs. In such cases, models must implicitly localize deviations based on high-level error cues (e.g., “action skipped”).
@@
Grauman et al. in Ego-Exo4D [grauman2024ego] explore a weak mistake detection strategy with two variants depending on supervision granularity: (1) instance-level supervision, where both video segments and key-step labels are provided during training and inference, analogous to action recognition, and (2) procedure-level supervision, where training and inference use unlabeled segments along with a taxonomy of procedure-specific key-step names. In the second approach, key-step assignments are generated via a pre-trained video-language model, providing pseudo-labels and exemplifying self-supervised weak supervision. Similarly, Ghoddoosian et al. [ATA] propose a weakly-supervised method for instructional mistake detection using only activity transcripts without frame-level labels or error examples. Supervision is limited to expected action sequences, with no temporal error information. Their method infers frame-wise action boundaries by aligning predictions to the transcript using a loss that enforces consistency with valid sequences while maximizing separation from invalid ones. Narasimhan et al. [narasimhan2023learning], in VideoTaskformer, operate under weak supervision differently: instead of aligning to a transcript, the model learns procedural knowledge by predicting masked steps within instructional sequences. Trained with step-level annotations (e.g., “crack egg”, “whisk”, “pour”), without step timing or correctness information, it infers missing steps based on surrounding context, acquiring a representation of procedural structure that identifies absent or out-of-order steps. In more recent works, Lee and Elhamifar [lee2025error] andJ Huang et al. [huang2025modeling] utilize structure-driven weakly supervised methods that leverage a predefined task graph as indirect supervision. Unlike transcript-alignment or masked-step prediction approaches, these methods do not require frame-level step labels or error annotations; instead, they perform joint step alignment and error recognition by optimizing a dynamic programming objective constrained by the graph topology. This framework allows both frame [lee2025error] and node [lee2025error, huang2025modeling] skipping, enabling the model to infer deviations from canonical execution without relying on densely annotated training data.
A complementary form of weak supervision is introduced by the PECC framework [ICME], which requires only step-level annotations of correct executions and no mistake labels. Unlike transcript-based alignment methods [ATA], masked-step prediction approaches [narasimhan2023learning], or graph-constrained frameworks [lee2025error, huang2025modeling], PECC does not rely on external procedural structure, textual cues, or predefined transition rules. Instead, it learns per-step feature distributions by training a temporal action segmentation backbone on correct sequences and fitting Gaussian Mixture Models in a one-class manner. Weak supervision here is defined purely through exposure to correct behavior, without auxiliary signals such as transcripts, task graphs, or LLM-generated pseudo-labels. PECC treats each step (action) independently, detecting errors as low-likelihood deviations in the learned feature space. This makes the approach effective for executional anomalies but limits its ability to capture higher-level procedural mistakes, such as step omissions or misordering, that other weakly supervised methods address through explicit global structure or sequence-level reasoning.
@@
Finally, a recent sub-line of weakly-supervised approaches [flaborea2024prego, plini2024ti] departs from previous formulations and defines mistake detection as a combination of step recognition and step anticipation. These methods exploit the structured nature of task transcripts as supervision signals, without requiring any frame-level action or mistake annotations. Notably, they leverage pre-trained LLMs as reasoning agents to infer expected next steps based on procedural understanding.